### v2

# Okapi: Efficiently Safeguarding Speculative Data Accesses in Sandboxed Environments

Abstract
This paper introduces Okapi, a new hardware/software cross-layer architecture designed to mitigate Transient Execution Side Channel (TES) attacks, including Spectre variants, in modern computing systems. Okapi enforces sandboxing for speculative execution, providing a hardware basis that can replace expensive speculation barriers in software.
At its core, Okapi allows for speculative data accesses to a memory page only after the page has been accessed non-speculatively at least once by the current trust domain. The granularity of the trust domains can be controlled in software to achieve different security and performance trade-offs. For environments with less stringent security needs, Okapi’s features can be deactivated to remove all performance overhead.
Without relying on any software modification, the Okapi hardware features already provide full protection against TES breakout attacks, e.g., by Spectre-PHT or Spectre-BTB, at a thread-level granularity. This incurs an average performance overhead of only 3.17 % for the SPEC CPU2017 benchmark suite.
Okapi introduces the OkapiReset instruction for additional software-level security support. This instruction allows for fine-grained sandboxing with custom program sizes smaller than a thread, resulting in 1.68 % performance overhead in our WebAssembly runtime experiment.
On top, Okapi provides the possibility to eliminate poisoning attacks. For the highest level of security, the OkapiLoad instruction prevents confidential data from being added to the trust domain after a sequential access, thereby enforcing weak speculative non-interference. In addition, we present a hardware extension that limits the exploitable code space for Spectre gadgets to well-defined sections of the program. Therefore, by ensuring the absence of gadgets in these sections, developers can tailor their software towards achieving beneficial trade-offs between the size of a trust domain and performance.
1 Introduction
Transient execution side channel (TES) attacks, including numerous variants of Spectre [28], persist as serious threats in modern computing systems. These vulnerabilities enable attackers to exploit side effects of speculative execution by tricking the processor into revealing information that is not accessed by the sequential semantics of the program. Since TES attacks exploit high-end features of modern microarchitectures such as speculative and out-of-order execution, eliminating them without introducing a significant performance penalty continues to be a major challenge. The most dangerous Spectre attacks involve so-called universal read gadgets that are capable of revealing the contents of any address in the entire memory space.
A widely used software mechanism to restrict a program’s accessible resources is sandboxing [44]. Untrusted code is often recommended to be executed within a sandbox [21] in order to prevent it from stealing data or otherwise compromising the system [36]. As long as no additional measures are in place, Spectre-style attacks pose a severe threat to such software-level isolation techniques [40]. Attackers can either perform a breakout attack, i.e., bypassing sandbox boundaries to access confidential data, or a poisoning attack, i.e., tricking a victim sandbox into transiently leaking its own memory content [34]. These attacks are particularly relevant to runtime environments that execute multiple (distrusting) tenants within a single thread. Securing runtime environments efficiently against breakout attacks is a major concern in industrial applications, e.g., serverless computing [40, 14].
Although microarchitectural patches have found adoption in commercial CPUs to mitigate Meltdown and MDS attacks, Spectre attacks are still primarily addressed by countermeasures at the operating system and compiler levels. These mitigations entail a high performance overhead [42, 2] but are used to patch legacy systems nonetheless. This is primarily because secure speculation techniques in hardware, to date, necessitate significant modifications to the microarchitecture and memory subsystem (additionally incurring substantial performance overheads) which impede their adoption in industrial settings.
1-A Contribution
We propose Okapi – a novel hardware/software cross-layer architecture for TES attack mitigation that provides breakout and poisoning security with low performance and hardware overheads. At its core, the Okapi architecture extends the Page Table Entries (PTEs) in the Translation Look-aside Buffer (TLB) with an additional bit, the safe access bit, that marks a data page whenever one of its addresses is translated legally, i.e., non-speculatively, without causing an exception, in the currently running thread. If this bit is set, the CPU is allowed to load speculatively from addresses that belong to this page. On the other hand, Okapi blocks speculative execution of loads from pages that do not have the safe access bit set and thereby prevents formation of any transient execution side channel based on the content of these pages. Whenever there is a switch between software threads, all safe access bits in the TLB are reset to prevent the misuse of stale access rights. With these hardware features in place, Okapi removes all universal read gadgets and provides security against breakout attacks at thread level.
In addition, Okapi introduces a dedicated ISA instruction, called OkapiReset, to enable the programmer to remove the access rights explicitly, e.g., after sensitive data regions have been accessed legally. This allows a designer to prevent breakout attacks at a finer granularity, e.g., at function level.
Besides breakout attacks, Okapi also mitigates poisoning attacks [35]. Okapi introduces efficient hardware features to drastically limit the exploitable code space for Spectre gadgets. Whenever the instruction pointer speculatively crosses instruction page boundaries, Okapi prevents speculative read accesses, which inhibits the initial phase of a Spectre-BTB [28] attack. This measure prevents exploiting any gadgets outside the current trust domain and restricts the set of potential gadgets for a poisoning attack to only those that are located on the currently executing code page. With this additional hardware mitigation in place, the software measures elaborated in Sec. 4-B offer a trade-off between manual effort and performance to efficiently eliminate poisoning attacks in a given application.
The main contributions of the paper are summarized as follows:
  • We introduce Okapi, a new hardware/software cross-layer architecture to effectively and efficiently protect against breakout and poisoning attacks (cf. security objective in Sec. 3). In particular, Okapi eliminates all threats imposed by Spectre-like universal read gadgets with only minor hardware and performance overhead.
  • Okapi offers a set of incremental security features. The hardware features incur low performance and implementation overhead and provide breakout security for thread-sized sandboxes. They drastically limit the set of exploitable gadgets for poisoning attacks. Additional software measures allow for breakout security at a custom granularity below thread-level and full poisoning security.
  • Okapi introduces a total hardware overhead of less than 0.4 % additional logic and flipflops (cf. Sec. 8). For use cases without security requirements, the Okapi features can be switched off to avoid all performance overhead. The Okapi hardware features, when used without any software modification, incur only a small performance overhead of 3.17 % (cf. Sec. 7), while significantly improving security at a thread-level granularity. Hence, Okapi blocks all sandbox breakout attacks, e.g., through Spectre-PHT and Spectre-BTB.
  • Further security measures at the software level involving the OkapiReset instruction enable Okapi to deliver breakout security in custom-sized sandboxes, with competitive performance. We provide a case study for protecting mutually distrusting applications in a WebAssembly runtime environment, introducing only 1.68 % performance overhead.
  • Security against poisoning attacks can be achieved by identifying and hardening the gadgets only within the code base of the victim sandbox. Alternatively, confidential data can be accessed using a special load instruction, OkapiLoad, which does not mark the accessed page as safe for subsequent speculative loads and therefore prevents any speculative access to that page. This, although at the cost of additional performance overhead, relieves the software developer from manually identifying gadgets within the victim’s code base.
The rest of the paper is organized as follows: Sec. 2 reviews background relevant for Okapi followed by Sec. 3 which formalizes the security objective. In Sec. 4 we introduce our Okapi concept and present its security implications in Sec. 5. Sec. 6 describes the Okapi architecture in detail before we evaluate different sandbox scenarios in a gem5 implementation of Okapi in Sec. 7. In Secs. 8 and 9, we analyze an RTL implementation of the Okapi hardware w.r.t. its hardware overhead and security guarantees. In addition, Sec. 9 inspects the size of trust domains in benchmark tests. Sec. 10 summarizes the state of the art of known mitigations for TES attacks.
2 Background
2-A Spectre Variants
TES attacks [11] exploit features of modern processors like speculative or out-of-order execution to transiently access sensitive data in memory that would not have been accessed in the sequential execution of the program. Every attack involves three main phases:
  1. 1.
    Setup: The adversary mistrains the targeted predictor and sets up the attacker process. Additional preparation like flushing the caches helps to enlarge the time window for the transient execution phase.
  2. 2.
    Transient Access + Transmit: The victim code snippet is executed and transiently accesses the secret memory region. A subsequent transient operation transmits the information by altering the microarchitectural state of the processor in a secret-dependent way. After the core recognizes the misprediction, it rolls back to the last committed state, while some microarchitectural footprints (e.g., the cache state) remain in the system.
  3. 3.
    Secret Recovery: The attacker recovers the secret data from the microarchitectural footprint and makes it visible in the architectural state, often using timing-based cache side channels like Flush+Reload [47] or a secret-dependent execution time in functional units [18].
The following part of this section briefly summarizes the three main classes of Spectre attacks:
1 cmp rbx, ARR_LEN
2 jg out-of-bounds // mistrained pred.: not taken
3 mov r8, [rax + rbx]
4 mov r9, [r10 + 4096 * r8] // leak secret
5 ...
6out-of-bounds:
7 ...
(a) Example pseudocode snippet of Spectre-PHT gadget
1 mov rbx, SECRET // load secret into rbx
2 jmp rax // speculate rax=leak
3 ...
4leak: // encode secret in cache state
5 mov rax, [B]
6 mov rcx, [rax + rbx]
(b) Example pseudocode snippet of Spectre-BTB gadget
1 mov [addr0], x // attacker controlled x
2 mov rax, [addr1] // speculate addr0==addr1
3 mov rdx, [rax] // load secret from x
4 mov rcx, [rbx + rdx] // leak secret
(c) Example pseudocode of Spectre-STL gadget
Figure 1: Spectre code gadgets (x86 pseudo-code)
Spectre-PHT
Also known as Spectre variant 1 [28], this attack exploits the pattern history table (PHT) that predicts the outcome of a branch decision. An example gadget is depicted in Fig. 1(a). During the setup phase, the adversary mistrains the predictor to predict the conditional ’jg’ jump to be not taken. Additionally, ARR_LEN is evicted from the cache to prolong the speculation time. Initiating the exploit, the attacker calls the code snippet with an out-of-bounds index rbx, targeting the memory address to be retrieved. Due to the mistraining, the processor assumes the provided index to be in bounds and executes both mov instructions. The gadget code transiently accesses the secret and stores it in the microarchitectural state of the data cache, by using the secret data as an address. After detecting the misprediction, the CPU reverts all program-visible effects of the transient instructions in the pipeline. However, the attacker can retrieve the data indirectly from the cache state. Because rbx is under the control of the attacker and can be of arbitrary value, the instruction sequence serves as a so-called universal read gadget.
Spectre-BTB/RSB
These variants, in contrast to Spectre-PHT, exploit speculation on indirect branches (variant 2 [28]) and return addresses (variant 5 [32, 29]). In this case, the predictors do not speculate on the outcome of a branch condition but on the target address the control flow jumps to. The attacker can poison the predictors to jump to any arbitrary gadget even within the victim’s code base to access the memory location of interest. Pseudocode for a Spectre v2 attack is depicted in Fig. 1(b). During the attack, the program counter makes a transient jump, leading to the execution of instructions that load secret data into the microarchitectural state before the misprediction is identified and corrected.
Spectre-STL
Besides control flow predictions that are used in Spectre variants 1 and 2, data flow predictions can open transient execution side channels as well ([1, 26]). Modern CPUs make predictions on target addresses within the load-store-unit to speed up data forwarding from one instruction to another. Not only does this pose a threat to data located in the store buffer, but it can also be used to construct a gadget that is able to transiently read from arbitrary memory addresses. An example of such a universal read gadget using data flow speculation is listed in Fig. 1(c). If the value of x is controlled by the attacker and the predictor is poisoned to predict the addresses addr0 and addr1 to match, the load-store-unit can forward the value of x to register rax. This value is then used as an address for the load instruction in line 3 to load from the attacker-controlled address x and afterwards transmit the content of this memory address through a microarchitectural side channel.
2-B Privilege Switches and Trust Domains
Processes
A process is an instance of a possibly multithreaded, computer program that is being executed. Each process has an execution context, consisting of the threads’ program counters, register contents, address space and privilege level.
Threads
Threads are independent sequential execution paths in a program that run concurrently or quasi-parallel. We say quasi-parallel here because it depends on the number of available (logical) cores of the underlying CPU hardware whether two threads can be executed in parallel or whether the operating system has to schedule them on a single core. The current hardware state of each thread consists of an instruction pointer and the register contents.
Privilege Levels
The privilege level of a process defines its rights to modify other processes. Typically, the operating system and device drivers run with a higher privilege level than ordinary applications. Processes with a higher privilege level are allowed to modify processes with a lower privilege level but not vice versa. Such modifications include changes in the address space of a process or even changes in register and memory values. An address space is a set of virtual addresses that are readable, writable, or executable by the threads of the process [41].
Privilege Switch
In the scope of this paper we understand a privilege switch as the change of the current privilege level under which the currently active thread on a CPU (logical) core executes. Privilege switches happen as part of context switches, when switching between threads, as well as upon system calls from user-level processes to privileged kernel code and their return back to user level. They are easily detectable in hardware.
Sandbox
A sandbox refers to the (predefined) address space (instructions and data) that a program is confined to by a trusted handler, e.g., the operating system or a runtime environment.
Trust Domains
In this paper, a trust domain denotes the dynamically growing set of data pages that the currently running program has, so far, accessed legally and that is therefore allowed to be accessed speculatively. Depending on the chosen granularity of the sandboxes (e.g., threads, jobs, functions, etc.), trust domains can grow to different sizes. When handing over control from one sandbox to another all access permissions are reset. This ensures that sandboxes are prevented from accessing each other’s data pages.
2-C Virtual Memory and Paging
Most modern computer systems employ virtual memory (VM) [5], providing the user with an abstract programming interface. The key concept is to use virtual addresses that are translated by the VM system to actual physical addresses to access the requested data. Virtual memory is organized in pages [^1] One page is typically 4 KB in size. Modern VM systems usually also support 2 MB or 1 GB pages which are often called hugepages or megapages. that are mapped to physical page frames of the same size. The mapping for each process is stored in a separate page table that is managed by the operating system. If a process accesses a virtual address, the VM system performs a page table walk to translate the virtual address to a physical address. Page table walks are considered to be slow as they involve several accesses to physical memory to look up the mapping. Additionally, VM annotates pages with permissions (read, write, execute), thereby implementing simple memory protection and isolation.
The VM subsystem is crucial for system performance since about one third of fetched instructions interact with the memory system [5]. Our own experiments with the SPEC CPU2017 benchmark suite, in which up to 25 % of instructions are loads, support this statement (cf. Sec. 7-A). Therefore, computer architects introduce dedicated hardware to speed up address translation. The Translation Look-aside Buffer (TLB) is a low-latency cache that keeps track of recent address translations to avoid costly page table walks.
3 Security Objective
The security objective of Okapi is to prevent leakage of speculatively accessed data residing in data memory or in the cache hierarchy, i.e., data at rest. Similar to existing hardware mitigations against transient execution side channels, like STT [48], Eager/Naïve Delay [37], and NDA [45], Okapi targets the objective of weak speculative non-interference [24].
Following the terminology of [24], we label our memory locations either as high (H), i.e., confidential, or low (L), i.e., public.
Breakout Security
We assume a program $p$ is vanilla-sandboxed, i.e., the high memory locations do not influence the architectural state under sequential program semantics. The program $p$ is breakout-secure or generally-sandboxed if it also satisfies weak speculative non-interference. This means that the high locations do not influence the architectural state under speculative semantics [24]. Provided that high and low locations do not share memory pages (i.e., vanilla sandboxes are implemented page-aligned), Okapi is guaranteed to enforce breakout security for $p$, as described in Sec. 4-A.
Poisoning Security
In contrast to breakout attacks, poisoning attacks target a victim program $p$ that has access to high memory locations. Additionally, $p$ has measures in place to protect high memory locations from being leaked sequentially. Nonetheless, an attacker can poison the control flow in $p$ to circumvent these measures under speculative semantics and leak confidential data [12, 34].
The victim program can be partitioned into two sets of sections $p_{H}$ and $p_{L}$. All sections of $p$ accessing high memory locations sequentially belong to $p_{H}$, while $p_{L}$ is the complement of $p_{H}$. In other words, ${p_{L}}$ is vanilla-sandboxed w.r.t. the high memory locations of $p$. The program must be partitioned in a way such that each individual section of $p_{H}$ cannot leak more than the data it accesses sequentially. Under such a partitioning, $p$ is poisoning-secure iff ${p_{L}}$ is breakout-secure w.r.t. the high memory locations. A trivial partitioning selects $p_{H}$ to contain only all load instructions to high locations. Sec. 4-B addresses the measures Okapi provides for poisoning security and discusses how to find an efficient and effective partitioning of $p$ into $p_{H}$ and $p_{L}$.
We consider all types of speculation that can open a transient window for a load instruction, including control-flow predictions, exceptions, memory-dependence speculation, and memory consistency violations.
Guarantees for speculative non-interference [24], i.e., ensuring that transient execution side channels do not reveal data in the register file and instruction operands (data in transit), are outside the scope of this paper. In addition, we exclude classical side channel attacks, e.g., monitoring the instruction cache footprint of square-and-multiply exponentiation in RSA [47]. Moreover, physical side channels, including power or electromagnetic side channels, fall outside the scope of this paper.
This security objective covers all TES attacks initialized by speculative loads, independent of the microarchitectural side channel being used to transmit and recover secrets (e.g., cache side channels [28], port contention [4], or data-dependent timing in functional units [3]). In particular, it covers the formidable Spectre-PHT and Spectre-BTB attacks for which, to date, no satisfactory mitigations exist in industrial practice.
4 Okapi Concept
Common to all Spectre variants targeting data at rest is a speculative load instruction that accesses a memory location that could not be accessed by the sequential program flow in that section of code. Preventing the execution of this transient access instruction is sufficient to avert an attack. Blocking all speculative loads from executing would achieve this goal but comes with a significant performance penalty, as shown in [27]. Most of the load instructions executed speculatively are not part of an attack. For minimum performance penalty, it is therefore crucial to only block the malicious ones.
Okapi builds upon this observation by blocking as few speculative load instructions as possible, taking into account legal accesses by the current trust domain. Our proposed architecture leverages the spatial and temporal locality in memory accesses, which means that most of the data retrieved within a short period of time is located in a narrow region in the memory space. Fig. 2 shows, for the SPEC CPU2017 rate suite, that programs indeed fulfill this locality assumption as they access only about 20 pages between context switches, on average. In fact, all benchmarks access at most 60 pages per context switch except for the first context switch, during which around 300 pages are accessed to load the benchmark binaries into memory.
In contrast to other isolation mechanisms, such as in-process isolation [43, 38], Okapi is not aware of a program’s address space and sandbox boundaries in advance and learns about legal accesses by tracking committed load instructions during the execution. Thus, a trust domain denotes the (dynamically growing) set of pages that the currently running thread has, so far, accessed legally and that can therefore also be accessed speculatively.
270270270270300300300300330330330330Total Code PagesCode Pages Mean per Context SwitchTotal Data PagesData Pages Mean per Context SwitchAverageperlbench_0perlbench_1gcc_0gcc_1gcc_2bwaves_0bwaves_1mcfcactuBSSNlbmxalancbmkdeepsjengimagickleelanabexchange2fotonik3droms00101010102020202030303030
Figure 2: The average number of accessed pages (code and data) in between context switches and the total number of allocated pages for the SPEC CPU2017 rate suite.
4-A Achieving Breakout Security
Okapi allows speculative access only to memory pages that have already once been accessed non-speculatively by the same sandbox. This confines potential transient attacks to only pages belonging to the current trust domain and removes all universal read gadgets capable of speculatively leaking data at arbitrary memory addresses by breaking out of the trust domain.
Okapi enforces speculative read permissions with a single safe access bit added to each page table entry in the data TLB. Whenever a page is accessed non-speculatively, the bit is set. Speculative loads are only allowed to obtain their physical address from a TLB entry if the corresponding safe access bit is set. Otherwise, the address translation is blocked until the load becomes non-speculative. This measure effectively prevents Spectre gadgets, such as the ones shown in Fig. 1, from accessing memory locations that do not have the safe access bit set and prevents breakout attacks for page-aligned sandboxes. In many cases, the first access to a page causes a miss in the TLB and subsequently triggers an expensive page table walk. Therefore, the delay induced by this mechanism in Okapi is largely amortized by the page table walk which reduces the overall performance overhead incurred.
Resetting the Permissions
During execution, the program architecturally accesses data pages and enables the corresponding safe access bits. Therefore, the trust domain, i.e., the amount of data that can possibly be leaked transiently, grows over time. Resetting the safe access bits at the right point in time is crucial to prevent the (mis-)use of stale access rights. The size of a program that establishes a trust domain can vary significantly, ranging from single functions to entire threads or processes. For correct handling of the access bits in Okapi, transitions between programs must be transparent to the CPU.
Sandboxes at process or thread level can easily be implemented in hardware. Privilege switches (cf. Sec. 2-B) are detected by monitoring the corresponding Control and Status Registers (CSRs) and automatically clear the safe access bits.
Therefore, as a first important result, we note that the Okapi hardware alone inhibits all universal read gadgets (e.g., the ones exploited by Spectre-PHT, -BTB and -STL, as depicted in Fig. 1) and achieves breakout security for thread-sized sandboxes.
In practice, it can be desirable to have more than one sandbox within the same thread or process to save costly context switches [15]. Detecting transitions from one sandbox to another below thread granularity is, unfortunately, not immediately possible at the hardware level. On the other hand, at the software level, sandboxes of any size are clearly distinguished. Therefore, Okapi empowers the software to communicate sandbox exits to the hardware.
OkapiReset Instruction
We add a dedicated instruction, OkapiReset, to the ISA that resets the trust domain by flushing all safe access bits. The OkapiReset instruction acts as a speculation barrier for load instructions. Similar to other speculation barriers, such as fences, it is executed only upon reaching the head of the Re-Order Buffer (ROB) to avoid premature resetting. As long as an OkapiReset instruction is present in the ROB, all younger load instructions are serialized to prevent out-of-order use of existing access permissions. Software developers can make use of the OkapiReset instruction to establish more fine-grained sandboxes, e.g., at function level.
In conclusion, the combination of the Okapi hardware and the OkapiReset instruction enforces breakout security for custom-sized sandboxes below thread level.
4-B Achieving Poisoning Security
While establishing trust domains for programs of a chosen granularity prevents transient accesses to the outside, data located on pages with an activated safe access bit can still be loaded speculatively. The challenge for Okapi w.r.t. such poisoning attacks (cf. Sec. 3) is to guarantee that those code sections $p_{L}$ within a program that do not sequentially access confidential data cannot misuse speculative access permissions gained by sensitive code sections $p_{H}$. We need to ensure that upon a transition from $p_{H}$ to $p_{L}$ the high memory locations are not contained in the trust domain. Placing confidential and public data on separate pages and augmenting the program with an OkapiReset instruction after every access to confidential data solves this problem as it effectively constrains $p_{H}$ to only contain all confidential load instructions. However, this incurs a substantial performance overhead as every OkapiReset instruction revokes all safe access bits, thereby also hindering speculation on previously accessed public data.
We therefore propose a second Okapi instruction which can preserve speculation on previously accessed public data. We introduce the OkapiLoad instruction that allows the software to access confidential data pages non-speculatively without adding the page to the trust domain, i.e., not setting the safe access bit. This ensures poisoning security as defined in Sec. 3 if every access to confidential data is performed by an OkapiLoad instruction.
Achieving full poisoning security for sandboxes that repeatedly process confidential data with these instructions comes at a high overhead due to serializing accesses to high memory locations. This can be relaxed by increasing the size of the sections in $p_{H}$ from single load instructions to larger segments of code to be able to speculate on high memory locations without leaking them speculatively.
A major challenge for increasing the size of the sections in $p_{H}$ are Spectre-BTB attacks, illustrated in Fig. 3. By poisoning branch prediction, an attacker can transiently divert the control flow from the sequential program flow (green arrow) to a gadget (blue arrow). Without additional measures, transiently executed gadgets can still access (red arrow) the content of pages already marked with a safe access bit, i.e., data pages accessed by $p_{H}$. This gadget can reside anywhere in the mapped code space, posing a challenge to conventional mitigations [46].
Therefore, we introduce an additional low-cost countermeasure directly in the hardware that drastically reduces the remaining gadget space. Okapi achieves this by delaying the execution of loads following a speculative control-flow decision crossing instruction page boundaries. Such loads can be detected in the fetch stage of the processor. If the next instruction to be fetched speculatively is located on a different page than the current instruction, the next instruction is flagged as suspicious (flag symbol in Fig. 3). During execution, every load instruction following a suspicious instruction is delayed until the control flow of the suspicious instruction becomes non-speculative. In Fig. 3 this is illustrated by Okapi intercepting the malicious read access. This measure efficiently prevents poisoning attacks that speculatively cross an instruction page boundary.
With this hardware measure in place, poisoning attacks are only possible if the control flow can (speculatively) reach a gadget without crossing page boundaries and the gadget targets confidential data on pages that are already a part of the trust domain, i.e., that have their safe access bit set. This facilitates the task for the software developer to protect against poisoning attacks at larger granularity (and, hence, lower performance cost).
We can enlarge the sections in $p_{H}$ as long as they fulfill the property of not containing any gadgets (cf. Sec. 3). Therefore, if we can guarantee a candidate region to be free of gadgets [46], we can extend the sections in $p_{H}$ to a chosen granularity, e.g., an entire function. The inspection of candidate regions can be done individually.
In conclusion, Okapi offers two different strategies to block poisoning attacks in sandboxes. The developer can either prevent confidential data pages from entering the trust domain by using the OkapiLoad instruction or revoke the access rights via an OkapiReset instruction when switching from $p_{H}$ to $p_{L}$. Note that already the hardware measures by themselves drastically reduce the code space available for exploitable gadgets as shown in our experiments in Sec. 9-B.
Applications that rarely access confidential data might benefit from guarding these accesses with OkapiLoad instructions and keeping the speculative access rights on public data for $p_{L}$. On the other hand, applications that repeatedly access confidential data benefit from larger code sections in $p_{H}$.
Trust Domain <Data Page> <Data Page> <Data Page> ... Secret data ... ... ... ... <Code Page> ... Jump to Func ... ... Func ... <Code Page> ... ... ... SpectreGadget ... ... poisoned branch!
Figure 3: Spectre-BTB challenge for Okapi. If speculative execution flow leaves the current page, all loads on the suspicious page are delayed until accesses to that code page are no longer speculative.
5 Security Reasoning
Okapi prevents speculative leakage of data outside a program’s memory space. The program counter of committed instructions as well as data and address of committed loads and stores of the program are considered exposed and therefore not protected. Compared to schemes that enforce weak speculative non-interference at an address granularity [48, 37, 45], Okapi achieves significantly better performance by declassifying an entire data page with the first non-speculative access to this page. As long as no OkapiLoad instruction is used, the Okapi hardware dynamically labels sequentially accessed data pages as low, i.e., public. This allows for a highly efficient protection against cross-sandbox breakout attacks solely in hardware.
As a result of the native hardware support for virtual memory translation based on page tables (cf. Sec. 2-C), most operating systems and runtime environments, such as in-process isolation [38, 35] and enclaves [7], utilize page granularity to define the accessible address space and enforce software boundaries. Commercial cloud services, such as Cloudflare workers [14], similarly incorporate page-aligned partitioning of the address space between mutually distrusting tenants. Therefore, mutually distrusting software domains typically possess non-overlapping page-aligned data memory address spaces, and page sharing only happens in instruction memory through shared libraries. Therefore, in the context of sandboxing, secure speculation with finer than page granularity has limited benefits while incurring a substantial performance penalty.
In the following, we discuss different scenarios for environments that benefit from the security guarantees provided by Okapi:
”Classic” Sandboxing
In classic sandbox scenarios where a trusted environment is protected from being accessed by an untrusted instance, we only need to ensure that all safe access bits are reset when handing over control to the untrusted job. If switching to the untrusted job involves a privilege switch, this is already taken care of by the Okapi hardware. Otherwise, an OkapiReset instruction needs to revoke all previously gained access rights.
Mutually Distrusting Sandboxes
Securing runtime environments that run multiple tenants within a single thread against breakout attacks is a major concern in industrial applications, e.g., serverless computing [40, 14]. Okapi secures this kind of sandboxed environments by preventing speculative cross-domain accesses even within the same virtual address space. Therefore, the runtime environment, e.g., the WebAssembly engine, needs only minor modifications to execute an OkapiReset instruction each time it enters and exits a sandbox, to fully separate the runtime and each sandbox from each other.
Vaulting
The inverse scenario to ”classic” sandboxing is vaulting, where a program is protected against malicious accesses from the outside. Besides breaking into the vault, attackers can launch a poisoning attack (cf. Sec. 3) to trick the victim program into speculatively leaking its own confidential data. Okapi already covers breakouts from other programs into vaulted code segments with the security mechanisms protecting mutually distrusting sandbox environments. In addition, Okapi mitigates poisoning attacks by employing the hardware measures described in Sec. 4-B. If desired, the OkapiLoad instruction prevents confidential data from being added to the vault’s trust domain.
6 Okapi Architecture
6-A Hardware Architecture
Trust DomainFetchDecodeRenameExecuteWritebackCommit1273ROB IQ 6 LSQ 45TLBsafe access bitInstruction Cache PortData Cache Port
Figure 4: Out-of-Order processor augmented with Okapi.
Okapi can be implemented in an out-of-order CPU as shown in Fig. 4. The modified parts are marked with numbers.
  1. Trust Domain1
    In the fetch stage every {PC, next PC} pair is compared to detect instructions that cross page boundaries. If the next PC is located on a different page the instruction is marked as suspicious for a possible Spectre-BTB/RSB attack.
  2. Trust Domain2
    The decode stage is extended to handle the OkapiReset and OkapiLoad instructions.
  3. Trust Domain3
    The ROB is modified to keep track of the visibility point, i.e., the youngest instruction in terms of program order that can open a transient window (cf. Sec. 3). Load instructions inside the ROB are marked as unsafe upon their insertion if they are younger than the visibility point and thus not certain to commit. An unsafe load instruction becomes safe, when it is passed by the visibility point, i.e., it is bound to commit. Similarly, loads are marked as suspicious_load if they are younger than an unsafe suspicious instruction. Loads between the visibility point and the first suspicious instruction clear their suspicious_load bit.
  4. Trust Domain4
    The Load Store Queue (LSQ) selectively blocks loads that are marked with the suspicious_load bit from going to the TLB and thus from executing speculatively.
  5. Trust Domain5
    Inside the TLB, safe loads behave as usual and set the safe access bit for the corresponding page upon a successful translation, while speculative and therefore unsafe loads can only be translated if the safe access bit for the corresponding TLB entry is already set. OkapiLoads behave analogously except for never setting the safe access bit. Privilege switches are detected by monitoring the CSRs storing the current privilege level. The safe access bits are cleared each time the privilege level changes. In addition, the TLB is modified to reset the safe access bits when the OkapiReset instruction is executed.
    Okapi eliminates Meltdown-style vulnerabilities by design since it only permits access to physical memory addresses in case of a valid, i.e., non-excepting, virtual address translation.
  6. Trust Domain6
    The Instruction Queue (IQ) of the processor is extended to reschedule all blocked load instructions. This includes the blocked suspicious_load instructions from the LSQ as well as unsafe loads that did not succeed in the TLB. All loads are rescheduled as soon as they are marked safe within the ROB while suspicious_load instructions can be rescheduled speculatively as soon as the suspicious_load bit is cleared.
  7. Trust Domain7
    The OkapiReset instruction is delayed until it reaches the ROB head, ensuring that all previous loads have been executed. The loads following the OkapiReset instruction are delayed until they become non-speculative or until the OkapiReset instruction is executed. This prevents the re-use of stale permissions.
The majority of the described extensions require only minor changes to the pipeline, like adding single bits to micro-operations or TLB entries. Okapi, to a great extent, re-uses the existing logic in the LSQ and requires no modification to the cache and memory subsystem. Implementing the Okapi instructions and the logic to determine the visibility point are the only significant hardware extensions.
6-B Additional Software Support
As described in Sec. 4, the Okapi hardware depends on additional software measures to effectively eliminate breakout attacks below thread level, and poisoning attacks. As a summary, the following measures taken at the software level are required to ensure the security guarantees:
  • Confidential data must not be placed on the same page as public, regularly accessed data. Thereby, only accesses to the confidential data pages need to be safeguarded and unnecessary slowdown in the non-sensitive part of the code is avoided.
  • Transitions from one sandbox below the size of a thread to another must perform an OkapiReset.
  • Transitions from critical code sections in $p_{H}$ to non-sensitive code sections in $p_{L}$ must perform an OkapiReset. This requirement can be relaxed if confidential data is only being accessed using OkapiLoad.
  • The sizes of the critical code sections in $p_{H}$ can be increased to, e.g., functions, if the instruction pages of these sections are known to not contain any gadgets.
The trusted entity, e.g., the operating system, needs to provide page-alignment for the data of each sandbox. The software developer can take care of the above measures by annotating secret data and using an extended compiler to insert OkapiReset and/or OkapiLoad instructions accordingly. In case the user wishes to increase the sections in $p_{H}$, the absence of gadgets can be accomplished by manual inspection or gadget analysis tools, e. g. [46, 22].
7 Performance Evaluation
The proposed architecture is implemented and evaluated based on the out-of-order CPU model o3 CPU provided by the gem5 simulator [6], configured according to Tab. I. First, we evaluate the performance of Okapi after implementing only the hardware features. This determines the overhead for enforcing breakout security in thread-sized sandboxes and limiting the gadget space for poisoning attacks to code pages that are accessed by the sequential data flow (Sec. 7-A). Next, we showcase how different jobs within the same thread of a WebAssembly runtime can easily be hardened against sandbox breakout attacks using the OkapiReset instruction (Sec. 7-B). Afterward, we examine the performance overhead caused by using the OkapiReset instruction to explicitly vault sensitive data in a cryptographic library (Sec. 7-C).
TABLE I: gem5 configuration
Processor
ISA x86_64
Decode width 5
Issue/Commit width 8
ROB entries 192
Instruction queue entries 64
LD/ST queue entries 32/32
Memory
DTLB entries 64
L1 I cache 32KiB, 8 ways
L1 I access latency 6 cycles roundtrip
L1 D cache 48KiB, 12 ways
L1 D access latency 6 cycles roundtrip
Private L2 cache 1280KiB, 20 ways
Private L2 access latency 60 cycles roundtrip
Main memory 16 GB DDR4 2400 MHz 8x8
7-A Thread-Level Sandboxing
Trust Domain2222333344445555Normalized execution timeover insecure baseline designNaïve DelayEager DelayDIFTOkapiGeomeanperlbench0perlbench1gcc0gcc1gcc2bwaves0bwaves1mcfcactuBSSNlbmxalancbmkdeepsjengimagickleelanabexchange2fotonik3droms0.80.80.80.811111.21.21.21.21.41.41.41.41.61.61.61.6
Figure 5: SPEC CPU2017 performance results
We benchmark the gem5 Okapi design by running the SPEC CPU2017 [8] reference workloads in full system mode using SimPoints [25]. The first 100 billion instructions are profiled to create a maximum of five SimPoints comprising 100 million measurement and 50 million warm-up instructions each for all workloads. The evaluation considers only the workloads that we were able to run on the unmodified o3 CPU model.
For comparison, besides Okapi, we implemented also other architectures that mitigate TES attacks targeting data at rest:
  • Naïve delay [37] – executes loads once they reach the head of the ROB.
  • Eager delay [37] – executes load instructions as soon as they are no longer speculative.
  • A dynamic information flow tracking (DIFT) architecture similar to [48, 31] – In DIFT, speculative load instructions are allowed to access the memory and initiate a taint for the accessed data. Speculative instructions that are able to form side channels are blocked from using tainted data.
For the experiments reported in this subsection neither OkapiReset nor OkapiLoad instructions were inserted into the benchmarks. The performance results for the SPEC CPU2017 suite are depicted in Fig. 5, summarizing the runtime overhead normalized to the insecure baseline design. It can be seen that Okapi outperforms all other mitigation schemes with an average overhead of only 3.17 %. The easy-to-implement blanket fixes, naïve and eager delay, incur a delay of roughly 2.6$\times$ or 2$\times$, respectively. Although it needs more complex modifications in the processor, the DIFT approach still imposes an overhead of 18.36 %. The significant performance increase originates in Okapi’s nature of adding an entire data page to the trust domain at a time. This means that it does not need to delay loads to the same page twice. The other examined fixes enforce weak speculative non-interference on address level. However, all architectures provide the same level of breakout security if the thread’s data is page-aligned.
[…]
Figure 6: Okapi: dispatched load instructions
[…]
Figure 7: Okapi: first-time issued load instructions
In the following, we analyze the results for the Okapi architecture in more detail. The vast majority of the load instructions enters the ROB as potentially malicious, as depicted in Fig. 6. On average, 85.5 % of the dispatched load instructions commit their results to the architectural state. However, this varies from 60.7 % in the case of mcf to 99.9 % for cactuBSSN. Fig. 7 illustrates the status of speculative load instructions when the LSQ first tries to execute their address translation. A load instruction can either be blocked due to being marked as suspicious_load or be passed to the TLB. On average, around 14.3 % of the speculative load instructions are blocked from going to the TLB due to our countermeasure to prevent the use of Spectre-BTB gadgets. Out of the remaining load instructions that are passed to the TLB speculatively, 98.4 % have their safe access bit set, confirming our locality assumption.
In general, workloads perform well if they have a high locality in combination with only few load instructions being blocked due to being marked as suspicious_load. This can be observed at the examples of bwaves and lbm. Delayed suspicious_load instructions do not always have a high impact on performance because they can be rescheduled speculatively as soon as the control flow speculation resolves. This can be observed at the example of perlbench_1 where 42.1 % of the speculative load instructions are not allowed to go to the TLB, initially. However, this workload still performs at the same speed as the insecure baseline.
On the other hand, in cactuBSSN, for example, 99.8 % of the dispatched load instructions commit, while only 91.3 % of the issued and unsafe load instructions have the safe access bit set. The delay by instructions that do not have the safe access bit set has a high impact on the performance since the speculation is correct in the majority of the cases, i.e., they commit later compared to the baseline design. This leads to the higher than average performance overhead of 6.5 %.
7-B Mutually Distrusting Sandboxes
We modified the Wasmtime WebAssembly runtime [10] to execute an OkapiReset instruction upon switching from one job to another. Resetting the permissions gained by the previously executed tenants effectively disables the possibility to mount a speculative breakout attack, even when multiple jobs are sharing the same thread.
In this experiment we split the Shootout workloads from the Sightglass Benchmark Suite [9] into two sets and reduce the iterations performed by the more computation intensive workloads to keep the simulation effort feasible. For each set, we set up the Wasmtime environment to execute the workloads interleaved. We run the two sets on our Okapi CPU as well as on an unmodified baseline CPU for comparison.
On average, Okapi incurs an overhead of only 1.68 % to mitigate breakout attacks between the different tenants.
7-C Vaulting
TABLE II: Libsodium Results
Config atomic $p_{H}$ function-sized $p_{H}^{\prime}$
runtime overhead #reset runtime overhead #reset
Hashing
sha256_64B 196.01 % 395 4.17 % 8
sha256_8KB 198.01 % 17794 4.67 % 135
Secret crypto
salsa20_stream 287.93 % 376 28.88 % 6
salsa20_stream_xor 171.66 % 440 7.29 % 6
secretbox_easy 170.69 % 1450 16.00 % 27
secretbox_open 182.16 % 1859 15.26 % 29
aead_ietf_encrypt 106.56 % 784 36.38 % 35
aead_ietf_decrypt 103.99 % 784 34.50 % 37
Public crypto
box_easy 154.54 % 214534 18.16 % 1324
box_open 155.04 % 214030 18.33 % 1313
Mean 168.43 % 45244.6 17.86 % 292
We benchmark the Libsodium [17] constant-time cryptographic library which already comes with a page-aligned memory allocator. This feature facilitates achieving poisoning security (cf. Sec. 6-B).
In this experiment we compare two different choices of granularity for $p_{H}$. First, we instrument the hashing, secret and public key routines within the cryptographic library to execute an OkapiReset instruction after every secret-dependent load instruction to prevent poisoning attacks from leaking secret data. This is the most conservative choice of $p_{H}$, also referred to as atomic $p_{H}$. Secondly, we adapt the same routines to reset the permissions on returns. This instrumentation approximates the security-critical parts of the program to the set of sections $p_{H}^{\prime}$ consisting of the code page(s) of entire functions. While this measure may not guarantee full security, it drastically reduces the attack surface as well as the potential gadget space in proportion to the region size (cf. Sec. 9-B). This can be automated w.r.t. to a predefined logical granularity. To this end, we modified the LLVM compiler framework [30]. LLVM’s intermediate representation allows for an easy adaptation of the chosen granularity. We refer to this instrumentation as function-sized $p_{H}^{\prime}$. The results in Tab. II serve as an estimate of the performance gain for the case that the developer can ensure the absence of gadgets in function-sized granularity, thus achieving full poisoning security.
Tab. II summarizes the runtime overhead of the median execution times of 100 runs of the different routines available in the library normalized to the unmodified version on an unmodified CPU. As can be seen, the finest possible granularity of $p_{H}$ introduces an average overhead of 168.43 %. Resorting to the more coarse-grained function-granular choice of $p_{H}^{\prime}$ significantly reduces the incurred overhead down to 17.86 %. Note that the Okapi security features against poisoning complement the security level provided by the constant-time library. Without Okapi, a poisoning attack can redirect the control flow such that the vault leaks the secret despite constant-time measures in place.
8 Hardware Overhead
We implemented the Okapi architecture according to Sec. 6 in the open-source BOOM processor. We ran FPGA synthesis for the mega configuration of BOOM that matches the performance of Intel Core 2 Duo in terms of CoreMark/MHz. In comparison to the baseline design, Okapi increases the number of lookup tables and flip-flops by 0.39 % and 0.35 %, respectively. This confirms our claim that the Okapi architecture incurs a negligible hardware overhead.
9 Security Evaluation
9-A Formal Verification
We base our verification on the methodology of Unique Program Execution Checking (UPEC), as presented in [19], which formally establishes weak speculative non-interference for a design under verification (DUV). Our DUV is BOOM extended by the described Okapi architecture. We adapt UPEC to sandboxing (cf. Sec. 3) by constraining the UPEC property [19] such that the program in its sequential semantics never accesses high memory locations.
The Okapi-adapted UPEC property holds after a proof runtime of 41 minutes. This guarantees breakout security for sandboxes of custom size, provided that the software executes an OkapiReset instruction on sandbox exits and that it places confidential and public data onto different data pages. In addition, this proves that any program $p$ is poisoning-secure if $p_{H}$ does not add the high memory locations to the trust domain.
9-B Attack Surface Analysis
As explained in Sec. 4-A, the number of speculatively accessible pages of a trust domain increases over time. However, the empirical results for SPEC CPU2017 shown in Fig. 2 demonstrate that programs access only 5.69 % of their mapped code pages and 31.27 % of their mapped data pages between two context switches [^2] Every context switch involves a privilege switch, on average. Thus, at most 31.27 % of a program’s data is accessible speculatively by any Spectre read gadget in the mapped address space. As Okapi prevents the execution of gadgets following transient jumps to arbitrary code pages, attackers could only exploit gadgets on the 5.69 % of the code pages that are reached by the program flow non-speculatively. This means that developers only need to care about Spectre gadgets in the immediate vicinity of locations (i.e., the pages) where the code works with secret data. Spectre gadgets in other locations are automatically suppressed by Okapi without any further action needed by the developers or the compiler.
An interesting observation we make is that, even if an attacker could successfully access confidential data, the cache side channel can only be exploited if the address used to encode the data is also within the trust domain. Although it is difficult to quantify the effect of this restriction on poisoning attacks, it is apparent that, without any software support, the Okapi hardware alone increases the necessary effort to exploit poisoning attacks drastically.
10 Related Work
SafeBet [23] exploits temporal locality based on similar observations as in Okapi. However, the SafeBet mitigation scheme introduces a cache-like hardware structure to keep track of which code regions have already accessed which memory regions legally. The performance overhead incurred by SafeBet depends on the chosen granularity of the monitored code regions, and ranges from 6 % (at 1 GB granularity) to over 270 % (at address level). Also, the hardware area overhead is significantly higher than that of Okapi, due to the additional 8.3 KB high-speed cache needed for access mapping.
Hardware-only mitigations based on DIFT, such as STT [48] or SecureBOOM [27], protect all data in the memory and caches. These schemes allow speculative accesses to data and track their effects through the pipeline with taint bits. Instructions capable of leaking their operands (“transmitters”) are then delayed until they are no longer speculative. The performance overheads of these mitigations range from 14.8 % to 36.0 %. DOLMA [31] and SPT [13] extend hardware-only mitigations to also protect the content of the register file (i.e., data that has been accessed non-speculatively) from being leaked speculatively while incurring a performance overhead of up to 45.0 %. Although these techniques (STT [48], SecureBOOM [27], DOLMA [31] and SPT [13]) deliver stronger security guarantees compared to the bare Okapi hardware, they come at the price of higher hardware and performance overheads. They significantly increase the design complexity of the microarchitecture, particularly across the memory hierarchy, which impedes their adoption in existing microarchitectures.
SpectreGuard [20] allows the programmer to annotate sensitive memory regions manually to protect data from Spectre-PHT attacks. The annotations are then used by the hardware to prevent memory accesses under control-flow speculation for all tagged regions by delaying their execution until they are safe to be executed. Protecting the content of the heap comes with 8.0 % performance overhead, while protecting the entire memory space used in the SPEC CPU2006 benchmark suite slows down the system by 20.0 %.
ConTeXT [39] pursues a related approach, where memory regions are marked as non-transient at a page-level granularity by the programmer, preventing access during transient execution. Additionally, ConTeXT employs DIFT to track sensitive data that has been loaded non-transiently into the register file, safeguarding such data from being used in transient operations. It requires intrusive changes over all layers (HW, OS, compilers) and incurs a baseline overhead of 1.26 % without protecting any data. With protection in place, the overhead reaches 71.14 % in security-critical applications.
ProSpeCT [16] also leverages the concept of manual annotations in software to extend secure speculation to constant-time programming, resulting in 10.0 % to 40.0 % performance overhead on custom benchmarks. Similar to ConTeXT, ProSpeCT is capable of protecting data at rest as well as data in the registers from TES attacks.
Okapi effectively enforces dynamic sandboxing for speculative data accesses and therefore shares this objective with conventional sandboxing approaches and (in-)process isolation techniques that also mitigate certain Spectre attacks [34, 38, 35, 7]. However, these mitigations pursue a different security target and introduce a much higher overhead in implementation and performance. In addition, while most of these techniques can thwart Spectre-PHT attacks, many of them leave an open attack surface to Spectre-BTB.
11 Conclusion
In this paper, we present Okapi, a novel hardware/software cross-layer architecture as an effective defense against breakout and poisoning attacks in sandboxed environments.
At its core, Okapi prevents these attacks by delaying speculative accesses to data pages which have not yet been accessed from within the currently running sandbox. Okapi provides the possibility to customize the granularity of sandboxes, thereby covering a range of possible applications. The Okapi hardware architecture is easy to implement and can be seamlessly integrated into existing designs. In particular, breakout security being highly relevant in applications of edge computing is achieved with an exceptionally low hardware and performance overhead. This bears promise for Okapi’s adoption in industrial practice, possibly replacing software patches deemed too costly.
An interesting direction for further research could be the combination of the Okapi hardware mechanisms with orthogonal software protection mechanisms, e.g., Serberus [33] for further performance improvements.
Acknowledgements
The reported research was partly supported by DFG SPP Nano Security, by BMBF grant SILKOSTU and by the Intel Corp. Scalable Assurance Program.